# The Critical LYZ equation in Kähler Geometry

### v1

@@ 4.4. Gradient estimates
Recall that $\hat{\theta}(t)\in(\pi-Ct,\pi)$, so $a_{0}\in[0,\infty]$. By passing to a further subsequence, we may assume
@@ 5. Examples: the $3$-dimensional case and $4$-dimensional case
As a corollary of our theorem, we solve the 2-Hessian equation under weaker condition instead of the usual condition $\chi\in\Gamma_{2}(M)$.
@@
Let $(M,\omega)$ be a compact Kähler manifold with dimension $n=3$ and $\chi$ be a real $(1,1)$ form satisfying the following conditions:
@@
    $3\int_{M}\chi^{2}\wedge\omega=\int_{M}\omega^{3}$, $\int_{M}\chi^{3}<3\int_{M}\chi\wedge\omega^{2}$, and
@@
    $\chi\wedge\omega>0\ \ \text{as a $(2,2)$ form}.$ (67)
@@
Condition $(i)$ implies the principal argument of $\int_{M}(\chi+\sqrt{-1}\omega)^{3}$ is $\pi$. Condition $(ii)$ means that 0 is the subsolution of the critical LYZ equation in dimension 3. According to Theorem 1.1, there exists a smooth solution solving the critical LYZ equation in dimension 3 which is exactly the 2-Hessian equation $\sigma_{2}=1$.∎
@@
If $\chi\in\Gamma_{2}(M)$, then it satisfies the subsolution condition $(ii)$. But the subsolution condition is weaker than $\chi\in\Gamma_{2}(M)$. We will prove the integral condition $(i)$ is necessary.
@@ 5.2. $4$-dimesional case
When $n=4$, the critical LYZ equation is the Hessian quotient equation. In this case the subsolution implies $\sigma_{2}(\mu|i)>1$ for $1\leq i\leq 4$, where $\mu=(\mu_{1},\mu_{2},\mu_{3},\mu_{4})$ are eigenvalues of $\chi$ with respect to $\omega$. Assuming further $\chi\in\Gamma_{3}(M)$, Sun [34] and Székelyhidi [36] proved the existence result using different continuity methods.
As a corollary of our result, we prove an existence result for the Hessian quotient equation $\sigma_{3}(\chi_{u})=\sigma_{1}(\chi_{u})$ under weaker conditions than before.
@@
Let $(M,\omega)$ be a compact Kähler manifold with dimension $n=4$. Let $\chi$ be a real $(1,1)$-form satisfying the following conditions:
@@
    $\int_{M}\chi^{3}\wedge\omega=\int_{M}\chi\wedge\omega^{3},\int_{M}\chi\wedge\omega^{3}>0,\ \mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{4}<0,$
@@
    $3\chi^{2}\wedge\omega-\omega^{3}>0\ \text{as a $(3,3)$-form}.$
@@
Condition $(i)$ implies the principal argument of $\int_{M}(\chi+\sqrt{-1}\omega)^{4}$ is $\pi$. To apply Theorem 1.1, we only need to show 0 is the subsolution of the critical LYZ equation. By condition $(ii)$, we have $\sigma_{2}(\chi)>2$ and then $\sigma_{1}^{2}(\chi)\geq\frac{8}{3}\sigma_{2}(\chi)>\frac{16}{3}$. Since there exists $z_{0}\in M$ such that $\sigma_{1}(\chi(z_{0}))=4\frac{\int_{M}\chi\wedge\omega^{3}}{\int_{M}\omega^{4}}>0$, we get $\sigma_{1}(\chi)>0$ on $M$.
Let $\mu=(\mu_{1},\mu_{2},\mu_{3},\mu_{4})$ be the eigenvalues of $\chi$ with respect to $\omega$ such that $\mu_{1}\geq\mu_{2}\geq\mu_{3}\geq\mu_{4}$. We need to prove
@@
In fact, by condition $(ii)$, we have $\sigma_{2}(\mu|1)=\mu_{2}\mu_{3}+\mu_{2}\mu_{4}+\mu_{3}\mu_{4}>1$. $\sigma_{2}(\mu|1)>0$ and $\mu\in\Gamma_{2}$ implies $\mu_{2}+\mu_{3}>0$ and then $\mu_{4}>\frac{1-\mu_{2}\mu_{3}}{\mu_{2}+\mu_{3}}$. Then we get
$\arctan\mu_{4}>\arctan\frac{1-\mu_{2}\mu_{3}}{\mu_{2}+\mu_{3}}=\frac{\pi}{2}-\arctan\mu_{2}-\arctan\mu_{3},$
which proves (72) and hence 0 is the subsolution of the critical LYZ equation. ∎
We next show condition $(i)$ is necessary.
@@
Let $u$ be the unique smooth solution of the Hessian quotient equation (71). Then $\chi$ satisfies condition $(i)$.
@@
The equality in condition $(i)$ follows from the integration of the equation. Since $\sigma_{1}(\chi_{u})>0$, we have
@@
Let $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})$ be the eigenvalues of $\chi_{u}$ with respect to $\omega$ such that $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}\geq\lambda_{4}$. By the MacLarin inequality, we have
@@
Since $\lambda_{4}\sigma_{1}(\lambda|4)=\lambda_{4}(\lambda_{1}+\lambda_{2}+\lambda_{3})\leq\lambda_{1}\lambda_{2}+\lambda_{2}\lambda_{3}+\lambda_{1}\lambda_{3}=\sigma_{2}(\lambda|4)$, we get
$\sigma_{2}=\sigma_{2}(\lambda|4)+\lambda_{4}\sigma_{1}(\lambda|1)\leq 2\sigma_{2}(\lambda|4),$ (75)
and then $\sigma_{2}(\lambda|4)\geq 3.$ Since $\lambda_{4}=\frac{\sigma_{3}(\lambda|4)-\sigma_{1}(\lambda|4)}{-\sigma_{2}(\lambda|4)+1}$, we get
@@
$=$ $\frac{\Big(\sigma_{3}(\lambda|4)-\sigma_{1}(\lambda|4)\Big)^{2}}{-\sigma_{2}(\lambda|4)+1}-\sigma_{2}(\lambda|4)+1$
@@
which proves the second inequality in condition $(i)$. ∎

### v2

@@ 4.4. Gradient estimates
Recall that $\hat{\theta}(t)\in(\pi-2Ct,\pi-Ct)$, so $a_{0}\in[0,\infty]$. By passing to a further subsequence, we may assume
@@ 5. Examples: the 3-dimensional case and 4-dimensional case
As a corollary of our theorem, we solve the 2-Hessian equation under weaker conditions instead of the usual condition $\chi\in\Gamma_{2}(M)$.
@@
Let $(M,\omega)$ be a compact Kähler manifold with dimension $n=3$ and $\chi$ be a real $(1,1)$-form satisfying the following conditions:
@@
    $3\int_{M}\chi^{2}\wedge\omega=\int_{M}\omega^{3}$, $\int_{M}\chi^{3}<3\int_{M}\chi\wedge\omega^{2}$;
@@
    $\chi\wedge\omega>0\ \ \text{as a $(2,2)$-form}.$ (68)
@@
Condition $(i)$ gives that the principal argument of $\int_{M}(\chi+\sqrt{-1}\omega)^{3}$ is $\pi$. Condition $(ii)$ means that 0 is a subsolution of the critical LYZ equation in dimension $3$. According to Theorem 1.1, there exists a smooth solution to the critical LYZ equation in dimension $3$, which is exactly the $2$-Hessian equation $\sigma_{2}(\chi_{u})=1$. ∎
@@
If $\chi\in\Gamma_{2}(M)$, then it satisfies the subsolution condition $(ii)$. However, the subsolution condition is strictly weaker than $\chi\in\Gamma_{2}(M)$. We will prove that the integral condition $(i)$ is necessary.
@@ 5.2. $4$-dimesional case
When $n=4$, the critical LYZ equation is the Hessian quotient equation. In this case, the subsolution condition yields $3\chi^{2}\wedge\omega-\omega^{3}>0$ as a $(3,3)$-form. Under the additional assumption that $\chi\in\Gamma_{3}(M)$, Sun [36] and Székelyhidi [38] proved the existence result using different continuity methods.
As a corollary of our result, we prove an existence result for the Hessian quotient equation $\sigma_{3}(\chi_{u})=\sigma_{1}(\chi_{u})$ under conditions weaker than those previously required.
@@
Let $(M,\omega)$ be a compact Kähler manifold of dimension $n=4$. Let $\chi$ be a real $(1,1)$-form satisfying the following conditions:
@@
    $\int_{M}\chi^{3}\wedge\omega=\int_{M}\chi\wedge\omega^{3}$, $\int_{M}\chi\wedge\omega^{3}>0$, and $\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{4}<0$;
@@
    $3\chi^{2}\wedge\omega-\omega^{3}>0$ as a $(3,3)$-form.
@@
Condition $(i)$ implies that the principal argument of $\int_{M}(\chi+\sqrt{-1}\omega)^{4}$ is $\pi$. To apply Theorem 1.1, we only need to show that 0 is a subsolution of the critical LYZ equation.
By condition $(ii)$, we have $\sigma_{2}(\chi)>2$, which implies $\sigma_{1}^{2}(\chi)\geq\frac{8}{3}\sigma_{2}(\chi)>\frac{16}{3}$. Moreover, since there exists a point $z_{0}\in M$ such that $\sigma_{1}(\chi(z_{0}))=4\frac{\int_{M}\chi\wedge\omega^{3}}{\int_{M}\omega^{4}}>0$, we conclude that $\sigma_{1}(\chi)>0$ on $M$.
Let $\mu=(\mu_{1},\mu_{2},\mu_{3},\mu_{4})$ be the eigenvalues of $\chi$ with respect to $\omega$, ordered such that $\mu_{1}\geq\mu_{2}\geq\mu_{3}\geq\mu_{4}$. We need to prove
@@
Indeed, condition $(ii)$ gives $\sigma_{2}(\mu|1)=\mu_{2}\mu_{3}+\mu_{2}\mu_{4}+\mu_{3}\mu_{4}>1$. Since $\sigma_{2}(\mu|1)>0$ and $\mu\in\Gamma_{2}$, it follows that $\mu_{2}+\mu_{3}>0$, and consequently
$\mu_{4}>\frac{1-\mu_{2}\mu_{3}}{\mu_{2}+\mu_{3}}.$
Therefore,
$\arctan\mu_{4}>\arctan\left(\frac{1-\mu_{2}\mu_{3}}{\mu_{2}+\mu_{3}}\right)=\frac{\pi}{2}-\arctan\mu_{2}-\arctan\mu_{3},$
which proves (73). Hence, 0 is a subsolution of the critical LYZ equation. ∎
We next show that condition $(i)$ is necessary.
@@
Let $u$ be the unique smooth solution to the Hessian quotient equation (72). Then $\chi$ satisfies condition $(i)$.
@@
The equality in condition $(i)$ follows directly from integrating the equation. Since $\sigma_{1}(\chi_{u})>0$, we have
@@
Let $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})$ be the eigenvalues of $\chi_{u}$ with respect to $\omega$, ordered such that $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}\geq\lambda_{4}$.
By the MacLaurin inequality, we obtain
@@
The inequality $\lambda_{4}\sigma_{1}(\lambda|4)\leq\sigma_{2}(\lambda|4)$ yields
$\sigma_{2}=\sigma_{2}(\lambda|4)+\lambda_{4}\sigma_{1}(\lambda|4)\leq 2\sigma_{2}(\lambda|4),$ (76)
and consequently,
$\sigma_{2}(\lambda|4)\geq 3.$ (77)
Using the identity $\lambda_{4}=\dfrac{\sigma_{3}(\lambda|4)-\sigma_{1}(\lambda|4)}{-\sigma_{2}(\lambda|4)+1}$, we compute
@@
$=\frac{\bigl(\sigma_{3}(\lambda|4)-\sigma_{1}(\lambda|4)\bigr)^{2}}{-\sigma_{2}(\lambda|4)+1}-\sigma_{2}(\lambda|4)+1$
@@
which establishes the second inequality in condition $(i)$. ∎